The electric motor of a model train accelerates the train from rest to 0.620 m/s in 21.0cs. The total mass of the train is 875 g. Find the average power delivered to the train during the acceleration.


1
Expert's answer
2021-11-07T19:24:25-0500

The average power delivered to the train during the acceleration can be found as follows:


P=Wt.P=\dfrac{W}{t}.

We can find the work done by the electric motor from the work - kinetic energy theorem:


W=KEfKEi=12mvf2.W=KE_f-KE_i=\dfrac{1}{2}mv_f^2.

Finally, we get:


P=0.875 kg×(0.620 ms)22×21.0×103 s=8.01 W.P=\dfrac{0.875\ kg\times(0.620\ \dfrac{m}{s})^2}{2\times21.0\times10^{-3}\ s}=8.01\ W.